OpenStudy (anonymous):

how do i solve the cubed root of 3xcubed ytothe 5?

OpenStudy (anonymous):

you can write that as cuberoot(xxxyyyyy); then take out the triples

OpenStudy (anonymous):

what about the cubed root of 24x^15 y^6?

OpenStudy (anonymous):

so take out 3 of the x and 3 of the y and that leaves two y, so \[xy \sqrt[3]{y^2}\]

OpenStudy (anonymous):

ok i had an answer similar to that, but not that exact one.

OpenStudy (anonymous):

take out as many triples as you can; and factor 24 into prime factors

OpenStudy (anonymous):

im not understanding the triples thing.

OpenStudy (anonymous):

24 is 2x2x2x3

OpenStudy (anonymous):

exactly... there are 3 2s so they can come out

OpenStudy (anonymous):

cuberoot(24) is 2 times cuberoot(3)

OpenStudy (anonymous):

fifteen xs so you can pull 5 triples out (all of them)

OpenStudy (anonymous):

so its 2x^5

OpenStudy (anonymous):

on the outside?

OpenStudy (anonymous):

or is the answer 2x^5 y^2?

OpenStudy (anonymous):

so far, but dont forget the ys

OpenStudy (anonymous):

yes... but there is something left under the radical

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you got it.... for square roots you pull out pairs, for cube roots you pull out the triples

OpenStudy (anonymous):

so the answer is 2x^5 y^2 cube root 3?

OpenStudy (anonymous):

I really appreciate your help. There is one more I need help with. It is under a square root, but is a fraction

OpenStudy (anonymous):

but be careful... this is the way to do single terms like these, but \[\sqrt{x^6 + y^2}\] is not x^3+y

OpenStudy (anonymous):

yes, that is the answer... dnag you are good and let's see the fraction

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

I'm not sure how to write it, but its cube root of 72x^15 y^13/64x^12 y^9

OpenStudy (anonymous):

\[\sqrt[3]{72x^15y^13 \over 64x^12y^9}\]

OpenStudy (anonymous):

simplify it first ... write the numbers as prime factors and cancel as much as you can

OpenStudy (anonymous):

i need some help with that. I don't understand.

OpenStudy (anonymous):

you can cancel 12 xs and 9 ys right off the bat

OpenStudy (anonymous):

so is it cube root of 72x^3 y^4/64xy?

OpenStudy (anonymous):

\[{x^{15} \over x^{12} } = x^3\]

OpenStudy (anonymous):

15 - 12 leaves 3 xs, and they happened to be on top

OpenStudy (anonymous):

so there are none left on the bottom, just the 64?

OpenStudy (anonymous):

exactly... got lucky this time

OpenStudy (anonymous):

with the problem I mean, you are doing great

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

so then i factor 72 and 64 or?

OpenStudy (anonymous):

now cancel out the common factors in 72 and 64

OpenStudy (anonymous):

yes, thats how to do it

OpenStudy (anonymous):

3x3x2x2x2?

OpenStudy (anonymous):

thats 72 it seems

OpenStudy (anonymous):

3x3x2x2x2 / 2x2x2x2x2x2 3x3 - - - - - - 2x2x2

OpenStudy (anonymous):

2x2x3x3?

OpenStudy (anonymous):

i know that last one is wrong.

OpenStudy (anonymous):

keep dividing by 2 until you can't do it any more then see if what is left is prime

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

64 is 2x2x2x2x2x2

OpenStudy (anonymous):

right so that is 3

OpenStudy (anonymous):

because there are 3 sets right?

OpenStudy (anonymous):

not yet, we are trying to simplify... cancelling what we can

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

that 72 is still up there, I mean the 3x3x2x2x2

OpenStudy (anonymous):

so 3x3x2x2x2 / 2x2x2x2x2x2 is 3x3 - - - - - - 2x2x2

OpenStudy (anonymous):

cacelled out 2x2x2 from top and bottom

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that leaves 9 x^3y^4 over 8 in the radical

OpenStudy (anonymous):

what goes on the outside?

OpenStudy (anonymous):

we havent pulled anything out yet, just simplified it you pick the triples to pull out (since this is a cube root)

OpenStudy (anonymous):

is it just one y on the outside?

OpenStudy (anonymous):

\[\sqrt[3]{9x^3y^4 \over 8}\]

OpenStudy (anonymous):

yes, pull 3ys out, one ouside and leaves one inside

OpenStudy (anonymous):

\[3y^{3}\] and \[x^{3}\] on the outside

OpenStudy (anonymous):

and then 9y/8 on the inside?

OpenStudy (anonymous):

almost... but when you take 3 of something from inside, it is 1 on the outside so an x and a y and the 1/8 can come out

OpenStudy (anonymous):

which does leave 9y inside ... and the cube root of 1/8 is 1/2

OpenStudy (anonymous):

1/8xy on the outside

OpenStudy (anonymous):

1/2 x y on the outside

OpenStudy (anonymous):

x is the cube root of x cubed .... y is the cube root of y cubed .... 1/2 is the cube root of 1/8

OpenStudy (anonymous):

so an x outside represents 3 inside, and so on

OpenStudy (anonymous):

ok so then its just x^2 y^2 on the inside?

OpenStudy (anonymous):

just a y on the inside?

OpenStudy (anonymous):

I think the answer is: \[{1 \over 2}xy \sqrt[3]{9y}\]

OpenStudy (anonymous):

you are right.

OpenStudy (anonymous):

you can check it by cubing what is on the left and multiplying that by what is under the radical

OpenStudy (anonymous):

there are lots of shortcuts, but they always confuse me I like doing things the long simple way

OpenStudy (anonymous):

okay great. thank you so much for your help. that was a hard one.

OpenStudy (anonymous):

only because of the typing... you knwo what you are doing keep up the good work

OpenStudy (anonymous):

thank you.